2.Cost-effectiveness Analysis

Amoria Phlebitis is a fictional illness from The Simpsons that causes sharp stabbing pains to the stomach and arms and temporary loss of vision. A 25-year-old patient diagnosed with this condition has an expected lifespan of 35 more years, during which they will continue to deal with these symptoms. Two new treatments (surgery and medication) have recently been proposed as alternatives to the current solution (no treatment).  

Medication prevents the disease from getting worse and will extend the patient's expected lifespan to 40 more years, and the pain and disability are reduced to a moderate level. The cost of medicine is estimated to be $12,000 for each additional year of life.

Surgery can completely cure the disease, but it also provides additional risks. The mortality rate of surgery is 4%. Of those who survive, it is estimated that 40% are fully cured of the disease. Doctors are able to tell if the patient is cured immediately after surgery. The cost of surgery is $150,000. If an individual is not cured, it is expected that they will immediately begin taking the medication for the rest of their life.

Surgery also comes with a 10% risk of serious infection. The infection is not life threatening, but it does increase expected costs by $100,000 and come with lifetime consequences, lowering a survivors utility values by 0.1 for each year of remaining life (regardless of whether or not the surgery was successful). All survivors have an expected lifespan of 40 years.

Individuals with Amoria Phlebitis have stated that, without treatment, each year of life is equivalent to 0.6 years in normal health (utility value of 0.6). The medication is expected to improve the value of each year with Amoria Phlebitis to 0.7. Individuals who are cured and do not get the severe infection are returned to excellent health (Utility value of 0.85).

A) Perform a cost-effectiveness analysis of these two treatments. This includes calculating the expected costs, life years, and QALYs from each of the three potential options.

B) Calculate all ICERs in terms of costs/QALY. If the government has an expected ICER threshold of $50,000 per QALY, what will the preferred treatment be?

C) Calculate the present value of the lifetime costs of Medication and Surgery using a 5% discount factor. Assume that the costs of surgery occur at the beginning of the year while the costs associated with an infection and medication occur at the end of the year.
Answer #1

If a person chooses to take medication, his expected life span increases to 40 years with a total cost of (40*12000)= $480000

If a person chooses to undergo surgery, then, with probability .04, he dies, and with probability 0.96 he survives. However, on survival, with probability 0.4, he is fully cured. Further , there is 0.1 probability he might be infected.

If the surgery is successful but he get infection, the cost is 150000(for surgery) and 100000( for infection), total $250000, with probability (0.96*0.4*0.1)=.0384, thus expected cost of $9600 with QALY (40*0.96*0.4*0.1*0.75)= 1.152

Without infection, a fully cured person has a cost of only $150000, with probability (0.96*0.4*0.9)=.3456, thus expected cost of $51840 with QALY (40*0.96*0.4*0.9*0.85)=11.75

If after surgery, patient is not cured and he gets infection, the cost is 150000(surgery), 100000(infection), 480000(medication), total $730000 with probability (0.96*0.6*0.1)=0.0576, thus expected cost of $42048, with QALY (40*0.96*0.6*0.1*0.5)= 1.152

If, after surgery, patient is not cured but he doesn't get infected, the cost is 150000(surgery) and 480000(medication), total $630000, with probability ( 0.96*0.6*0.9)=0.5184, thus expected cost of $326592, with QALY (40*0.96*0.6*0.9*0.7)= 14.52

Therefore, total expected cost of surgery is $430080.

Without any treatment, QALY is 21(0.6*35)

With medication, QALY is 28(0.7*40)

With surgery, expected QALY is 28.574(1.152+11.75+1.152+14.52)

B.  

If Government has a threshold of $50000 per QALY, the preferred treatment would be surgery as its expected cost is cheaper than 40 years' medication and the highest cost per QALY for a person with no treatment is (430080/21)= 20480 which is lower than 50000.

C.

The present value of lifetime cost of medication of $12000 per year for 40 years is 12000*[{1-1/(1+.05)^40}/.05] equals approximately $205909

The present value of surgery is $150000 incurred at the start of 1st year

Present value of cost of infection is $105000, incurred at thetend of year 1.

> where did you get the 0.75?

Elaine San Agustin Sat, Nov 20, 2021 7:49 PM